Intro Graphics April 9, 2006 G. Farin

Programming Assignment # 4 due: 4-26 percent of your your grade: 12.5

For this project, the idea is to add textures to objects such as those you used in the previous project.

Implement: bump maps texture files (read in .bmp file and apply it) Toon shading via 1D textures

Use an object which is composed of rectangular patches, such as the Utah teapot.